Can someone help me solve this? Thank you.
Greg bought a desktop computer and a laptop computer. Before finance charges, the laptop cost $200 less than the desktop. He paid dot the computer using two different financing plans. For the desktop the interest rate was 8% per year, and for the laptop it was 6.5% per year. The total finance charges for one year were $335. How much did each computer cost before finance charges?

Answer by stanbon(75887)

Greg bought a desktop computer and a laptop computer.
Before finance charges, the laptop cost $200 less than the desktop.
Let desktop cost be x
Then laptop cost is x-200
-------------------------------
He paid for the computers using two different financing plans.
For the desktop the interest rate was 8% per year, and for the laptop it was 6.5% per year.
Interest on desktop = 0.08x
Interest on laptop = 0.065(x-200)
------------------------------
The total finance charges for one year were $335.
How much did each computer cost before finance charges?
======
Equation:
interest + interest = interest
0.08x + 0.065(x-200 = 335
Multiply thru by 1000 to get:
80x + 65x - 65*200 = 335000
----
145x = 348000
x = $2400 (cost of the desktop)
x-220 = $2200 (cost of the laptop)
======================================
Cheers,
Stan H.
======================================